ICM 1998, Berlin, Aug. 18–27

Abstracts of Short Communications and Posters

Section:

8. Analysis

1991 MS Classification: 30F10 Hidalgo, Ruben A., UTFSM, Chile A_{4}, A_{5} and S_{4} of Schottky Type Let H be a group of conformal automorphisms of a closed Riemannn surface S. Assume that H is isomorphic to either the alternating groups A_{4} or A_{5} or to the symmetric group S_{4}. We provide necessary and sufficient conditions for H to have the existence of a Schottky uniformization of S for which H lifts.